A bag contains 3 white balls 4 black balls and 5 red balls if 3 balls are drawn at random determine the probability that.
1) all three red balls
2) all three are black
3) 2 are red and 1 is black
4) one of each Color is drawn
5) one is white
6) at least 1 white
1
Expert's answer
2020-04-30T19:17:50-0400

Solution (1): The total number of outcomes when pulling a single ball (no matter what) is equal to the sum of all types of balls "N_1=3+4+5=12". The total number of outcomes when pulling a red ball is the number of red balls "n_{r1}=5" . The probability to get red ball first is "P_{r1}=n_{r1}\/N_1=5\/12" . Before the second attempt of pulling out a red ball, the situation has changed according to

"N_2=3+4+4=11" , "n_{r2}=4" . Thus "P_{r2}=n_{r2}\/N_2=4\/11" is a probability to get red ball in a second attempt. The probability of getting a red ball in the third attempt is also calculated "P_{r3}=3\/10". The total probability is a product of the probabilities of successful attempts:

"P_{(1)}=P_{r1}\cdot P_{r2}\cdot P_{r3}=\frac {5\cdot 4\cdot 3}{12\cdot 11 \cdot 10}=1/22"

Solution (2): Reasoning similarly to point (1) we get

"P_{(2)}=P_{b1}\cdot P_{b2}\cdot P_{b3}=\frac{4\cdot 3\cdot 2}{12\cdot 11\cdot 10}=1/55"

Solution (3): The total number of outcomes when pulling three balls out of 12 is equal to number of combinations "(^{12}_3)"which is convenient to write here as "C_{12}^3" [1] "N=C^3_{12}=\frac{12!}{9!\cdot 3!}=220"

The total number of ways to pull two red balls out of 5 lying in the bag is equal to "n_{2r}=C_5^2" . The total number of ways to pull one black ball out of 4 lying in the bag is equal to "n_{1b}=C_4^1=\frac {4!}{1!\cdot 3!}=4" It was and just so clear. The probability of a positive result is estimated as a quotient of dividing the number of positive outcomes by the total number of cases

"P_{(3)}=\frac{n_{1b}\cdot n_{2r}}{N}=\frac{4\cdot C_5^2}{C_{12}^3}=\frac{4\cdot 5!/(2!\cdot 3!)}{12!/(3!\cdot 9!)}=\frac{4\cdot 5! \cdot 3!\cdot 9! }{12!\cdot 2!\cdot 3!}=2/11"

Solution (4): Using the results of the previous arguments we get "n_{1r}=C_5^1=5, n_{1b}=4, n_{1w}=3" and "P_{(4)}=\frac{n_{1r}\cdot n_{1b}\cdot n_{1w}}{N}=\frac{5\cdot 4\cdot 3}{12!/(3!\cdot 9!)}=\frac{5\cdot 4\cdot 3\cdot 3!\cdot 9!}{12!}=" 3/11

Solution (5): The total number of outcomes when pulling two balls not white is "n_{2rb}=C_{(12-3)}^2=C_9^2" Thus "P_{(5)}=\frac{n_{1w}\cdot n_{2rb}}{N}=\frac{3\cdot C_9^2}{C_{12}^3}=\frac{3\cdot 9!/(2!\cdot 7!)}{12!/(3!\cdot 9!)}=\frac{3\cdot 9!\cdot 9!\cdot 3!}{2!\cdot 7! \cdot 12!}=27/55"

Solution (6): The number of outcomes when there are no white balls in the sample is equal to"n_{3rb}=C_{(12-3)}^3=C_9^3" . To get the number of outcomes with any number of white balls, you need to subtract the obtained value "n_{3rb}" from the total number of outcomes "N" from previous. That is

"n_{w}=N-n_{3rb}" . Finally we get

"P_{(6)}=\frac{n_w}{N}=1-\frac{n_{3rb}}{N}=1-\frac{C_9^3}{C_{12}^3}=1-\frac{9!/(3!\cdot 6!)}{12!/(3!\cdot 9!)}=1-\frac{9!\cdot 9!}{12!\cdot 6!}=1-21/55=34/55"

Answers:

1) "\frac{1}{22}" ; 2) "\frac{1}{55}" ; 3) "\frac{2}{11}" ; 4) "\frac{3}{11}" ; 5) "\frac{27}{55}" ; 6) "\frac{34}{55}"


Note: "P_{(1)}=\frac{C_5^3}{C_{12}^3}; P_{(2)}=\frac{C_4^3}{C_{12}^3}" ; "C_n^m=(^{n}_m)=\frac{n!}{m!\cdot (n-m)!}"

[1] https://en.wikipedia.org/wiki/Combination